Find three consecutive integers so that the square of the second increased by the product of the other two is 199.
Answer by stanbon(75887)

Find three consecutive integers so that the square of the second increased by the product of the other two is 199.
-----
1st: x-1
2nd: x
3rd: x+1
-----
Equations:
x^2 + (x^2-1) = 199
2x^2 -200 = 0
2(x^2-100) = 0
x^2-100 = 0
(x-10)(x+10) = 0
x = 10 or x = -10
-----
If x = 10,
1st: x-1 = 9
2nd: x = 10
3rd: x+1 = 11
----
If x = -10,
1st: x-1 = -11
2nd: x = -10
3rd: x+1 = -9
==================
Cheers,
Stan H.
==================